∠A and \angle B∠B are complementary angles. If m\angle A=(x-21)^{\circ}∠A=(x−21)

and m\angle B=(2x-15)^{\circ}∠B=(2x−15)

, then find the measure of \angle A∠A.

Answer by ikleyn(52778)

.
∠A and \angle B∠B are complementary angles. If m\angle A=(x-21)^{\circ}∠A=(x−21)

and m\angle B=(2x-15)^{\circ}∠B=(2x−15)

, then find the measure of \angle A∠A.
~~~~~~~~~~~~~~~~~~

Your equation is


    m(A)   + m(B)    = 90 degrees    (complementary angles)

or
    
    (x-21) + (2x-15) = 90.


Simplify and find x


    x - 21 + 2x - 15 = 90

    x      + 2x      = 90 + 21 + 15

         3x          = 126

          x          = 126/3 = 42.


ANSWER.  m(A) = x-21 = 42-21 = 21  degrees;   m(B) = 2x-15 = 2*42 - 15 = 69 degrees.


CHECK.   m(A) + m(B) = 21 + 69 = 90  degrees.     ! Correct !

Solved.

-------------------

To learn about complementary angles,  read the lessons
    - Angles basics
    - Vertical angles
    - HOW TO solve problems on supplementary, complementary or vertical angles - Examples
    - Solved problems on supplementary and complementary angles
in this site.


Consider these lessons as your handbook,  textbook,  guide,  tutorials and  (free of charge)  home teacher.